can you please help i forgot how to do this
Solve each equation by the subtraction method. Determine whether the equation are independent, dependant, or inconsistent.
28. x =y + 4
3y – 2y = 4
thank you in advance

Answer by funmath(2933)

Here you go:
L1) x=y+4
L2) 3x-2y=4
---------------Since L1 is already solved for x, substitute that in for x in L2 and solve for y
3(y+4)-2y=4
3y+12-2y=4
(3-2)y+12=4
1y+12=4
y+12=4
y+12-12=4-12
y=-8
---------------Substitute -8 into L1 for y and find out what x is:
x=(-8)+4
x=-8+4
x=-4
The solution is: (x,y)=(-4,-8)
:
Check by substituting this solution into both equations and see if they equal out.
L1) -4=-8+4
-4=-4 L1 Checks out.
:
L2) 3(-4)-2(-8)=4
-12+16=4
4=4 L2 Checks out also.
:
Now we can be sure that we were right, the solution really is, (-4,-8)
Happy Calculating!!!